Disaster Cleanup stock sells for $29 a share. The firm has a rights offer outstanding for new equity shares. Antonio currently owns 400 shares. He just received one right for every share he owns. To purchase one new share he must submit five rights and $22. What is the value of one right?

Multiple Choice

  • $1.17

  • $1.33

  • $.91

  • $.84

  • $1.09
Answer #1

Value of one Right = (Stock Price - Rights subscription price per share) / (Number of rights required to buy one share + 1)

Stock Price = $29

Rights subscription price per share = $22

Number of rights required to buy one share =5

Value of one Right = (29-22) / (5+1) = 7/6 = 1.17

answer is a